f(x)=-4x^2+20x
Use the vertex formula to find the vertex.

Answer by John10(297)

Apply the vertex formula: V = (-b/2a , f(-b/2a))
x = -b/2a = -20/2(-4) = -20/-8 = 5/2
y = f(5/2)= -4(5/2)^2 + 20(5/2) = -4(25/4) + 50 = -25 + 50 = 25
Thus the vertex is (5/2, 25)
John10:)